a rectangle is 12 meters longer than it is wide. its perimeter is 68 meters find its length and width.
Answer by mananth(16946)

let width be x
length will be x+12
..
Perimeter = 2*(L+W)
2*(x+x+12)=68
2*(2x+12)=68
4x+24=68
add -24
4x+24-24=68-24
4x=44
x= 11 m the width
length = 23 m
CHECK
23+23+11+11==68